512 Hydraulically smooth pipe walls. water av -51/s 2m
Answer #1

flowrate qv = 5 L/s = 0.005 m3/s

velocity inside the pipe v = qv / Ac

Ac = cross-sectional area of the pipe.

d = 50 mm = 0.05 m

v = \frac{q_{v}}{\frac{\pi }{4}d^{2}}

v = \frac{0.005}{\frac{\pi }{4}0.05^{2}} = 2.54648 \ m/s

density of water = 1000 kg/m3

\nu = \frac{\mu }{\rho }

\mu = 1000 x 1.3 x 10-4 = 0.13 kg/m.s

Re=\frac{\rho vd}{\mu } = \frac{1000 *2.54648*0.05}{0.13} = 979.415

As Re < 2100, flow is laminar.

For laminar flow, the frictional head loss is given by

h_{L} = \frac{32\mu vL}{\rho gd^{2}}

h_{L} = \frac{32*0.13*2.54648*10}{1000*9.81*0.05^{2}} = 4.3194 \ m

Bernoulli's equation for point 1 and 2,

\frac{P1}{\rho g}+\frac{v1^{2}}{2g}+z1 = \frac{P2}{\rho g}+\frac{v2^{2}}{2g}+z2 + h_{L}

P2 = Po

v1 = v2 \approx 0

z1 = 1.2 m

z2 = 0.8 m

\frac{P_{1}}{1000*9.81}+0+1.2 = \frac{P_{0}}{1000*9.81}+0+0.8 + 4.3194

.: P1 - P0 = 38449.314 Pa...............................................................Answer